7x + 2y =7
2x + y = -1
1
Expert's answer
2011-11-15T09:36:40-0500
7x+2y=7
4x+2y=-2

substract the second equation from the first:
3x=9 => x=3
from the second equation of the system 2y=-2-4x=-2-12=-14 => y=-7
Hence answer x=3, y=-7